OpenStudy (anonymous):

plz help: the expression sin(theata+90) is equivalent to?

OpenStudy (cwrw238):

cos theta

OpenStudy (anonymous):

how do u no tho?

OpenStudy (anonymous):

@AccessDenied canu plz help me

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@Koikkara

OpenStudy (anonymous):

@Kilorion

OpenStudy (anonymous):

@hoblos

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

canu plz explain it more?

OpenStudy (loser66):

no I m sorry, it's cos theta

OpenStudy (anonymous):

canu help me with another ques plz?

OpenStudy (loser66):

sin (a+ b) = sina cos b + sinb cos a your a = 90 , your b = theta so that sin (90 + theta) = sin 90 cos theta + sin theta cos 90 sin 90 =1, cos 90 = 0 so that sin (90+ theta ) = 1* cos theta + sin theta *0= cos theta

OpenStudy (loser66):

I have class right now, cannot help you more. I am sorry, but I can give you the answer if the problem is easy, hurry up

OpenStudy (anonymous):

: the expression cos4x cos3x+sin4x sin3x is equivalent to?

OpenStudy (loser66):

cos x because cos (a-b) = cos a cos b + sina sinb your a is 4x, your b is 3x, so that cos (4x-3x) = cos x

OpenStudy (loser66):

gtg, good luck

OpenStudy (anonymous):

thnks so much ur awesome